# FuSeFL: Fully Secure and Scalable Cross-Silo Federated Learning

### v1

@@ Abstract
1 Introduction
Federated Learning (FL) is a distributed machine learning paradigm that enables multiple clients to collaboratively train a global model without sharing their local data [1, 2, 3, 4, 5, 6, 7]. This paradigm is particularly suited for scenarios where legal, privacy, or infrastructural constraints prevent centralized data collection [8, 5]. In FL, a central server distributes a model to the clients, who perform local training and return updates, such as gradients or model weights. These updates are aggregated to refine the global model, all while the raw data remains on client devices, preserving confidentiality and reducing transmission costs.
While FL offers clear advantages in preserving data locality, it remains vulnerable to indirect privacy breaches. In particular, submitted gradients or model updates can unintentionally encode patterns from the underlying data, enabling adversaries to extract sensitive information from these values. For example, attackers may perform model inversion to reconstruct private inputs or launch membership inference attacks to determine whether a particular data point was part of the training set [9].
To mitigate such risks, a wide range of techniques have been proposed to protect the privacy of clients’ data during training. Homomorphic encryption (HE) enables secure aggregation over encrypted updates without decryption [3, 4], but incurs high computational and memory costs, often slowing down training by several orders of magnitude [10]. Differential privacy (DP) defends against inference attacks by adding noise to updates [5, 6], though stronger privacy budgets degrade model accuracy [10]. Secure multi-party computation (MPC) allows collaborative aggregation without exposing individual inputs [7, 11, 12, 13], but suffers from growing communication overhead as the number of clients increases [10].
Beyond data privacy, a critical yet often underappreciated concern in federated learning is the privacy of the model itself, particularly from the perspective of the model owner. In typical FL deployments, the global model, often developed through significant intellectual and financial investment, is shared with all participating clients, some of whom may be semi-honest or untrusted [14]. This exposes the model to risks such as extraction, reverse engineering, or unauthorized use. The concern becomes even more pressing in cross-silo FL, where clients are typically large institutions, such as hospitals, banks, or research labs, that operate under strict privacy regulations and contractual obligations. In such settings, the model is not only a valuable proprietary asset but also a potential vector for secondary leakage of sensitive information. Ensuring model confidentiality is therefore essential, not only to protect intellectual property but also to comply with legal, regulatory, and contractual boundaries.
@@ 1 Introduction
WW-FL [17] adopts a fully offloaded training architecture in which secure multi-party computation (MPC) is performed across distributed clusters of servers. While this design offers strong security guarantees, it requires clients to upload their entire datasets to third-party servers, contradicting the foundational principle of cross-silo FL, where data should remain local to each silo. This architecture introduces significant communication overhead, as large volumes of sensitive data must be transmitted over the network, increasing both latency and memory usage. Moreover, it depends on geographically distributed compute infrastructure, which elevates deployment complexity and operational costs. In cross-silo environments, where clients such as hospitals or financial institutions already possess rich compute and storage capabilities, outsourcing training to external servers is not only inefficient but also unnecessary.
@@
AriaNN [22] introduces a hybrid federated scheme that jointly protects model and data privacy by secret-sharing the model between the client and server and distributing the computation using MPC. This design prevents either party from observing the full model or training data in plaintext. While more aligned with FL principles, AriaNN suffers from serious scalability limitations. Each client must train its share of the model in coordination with the server, typically in a sequential fashion due to server-side compute constraints. Communication overhead grows linearly with client count; even modest datasets like MNIST can require tens of megabytes per client, and this cost becomes prohibitive for large datasets common in siloed environments. Furthermore, the server must maintain a dedicated model instance per client in addition to the global model, leading to unsustainable memory overhead for large-scale models such as VGG-16.
@@
In FuSeFL, clients are dynamically paired to form collaborative groups. Within each group, the global model is secret-shared between the two clients, and training is jointly executed using secure multi-party computation (MPC) over their local data. This structure prevents any party from accessing the model or data in plaintext. After training, each group submits secret-shared updates to two non-colluding aggregation servers, which perform secure aggregation without learning individual contributions.
@@
  • A fully secure FL scheme: We develop FuSeFL, a federated learning scheme that holistically secures model parameters, client data, and intermediate updates using lightweight MPC across client groups—addressing both model and data privacy within a unified design (§5.1).
  • Decentralized and scalable training: FuSeFL eliminates server-side training bottlenecks by distributing computation among client pairs. This design supports parallelism, avoids reliance on third-party compute infrastructure, and is well-suited for cross-silo FL environments with powerful clients (§7.2).
  • Practical efficiency with robust privacy guarantees: Our approach significantly reduces memory and communication overheads while mitigating key inference threats such as model inversion, gradient leakage, and membership inference, making FuSeFL scalable to real-world deployments involving many silos (§7.4, §7.3).
2 Motivation and Goal
2.1 Limitations of Existing Work
@@
    Model Exposure and Trust Assumptions: Many existing FL schemes expose the full global model to clients during training or inference, risking model theft or misuse, particularly in regulated domains where model intellectual property is sensitive [25, 26, 15, 16]. Moreover, several privacy-preserving methods offload client data or models to trusted third-party servers using MPC or HE protocols [19, 20, 21, 22, 27, 28]. These trust assumptions violate the core principle of cross-silo FL, which requires data to remain within institutional boundaries [29, 30].
@@
    Communication and Storage Overheads: Secure training introduces high communication and memory demands. For example, training Network1 over MNIST with AriaNN-FL requires more than 10 MB of data exchange per client per epoch [22]. With hundreds of clients, this quickly overwhelms bandwidth limits. Additionally, the server must retain a copy of each client’s model update until aggregation is complete. For large models like VGG-16, this results in excessive memory usage: as illustrated in Fig. 1 (right axis), AriaNN-FL stores 128 full model instances when training with 128 clients.
@@ 2.2 Our Goals
3 Threat Model and Security Considerations
@@
These guarantees mitigate risks such as unauthorized observation, gradient leakage, membership inference, model inversion, and reconstruction attacks. FuSeFL enforces these protections using efficient secure computation protocols based on additive and function secret sharing.
@@
4 Background
4.1 Notations and Preliminaries
Our scheme leverages MPC to preserve the privacy of unencrypted data during training. Table I summarizes the notations used throughout the paper.
@@ 4.1 Notations and Preliminaries
$GM$ Global model in plaintext
$[GM]_{j}$ $j^{th}$ share of the secret shared global model
@@
$[LM_{i}]_{j}$ $j^{th}$ share of the secret shared $i^{th}$ local model
@@
$[D_{i}]_{j}$ $j^{th}$ share of the secret shared data of $i^{th}$ client
4.2 Federated Learning (FL)
@@
4.3 Secure Multi-Party Computation (MPC)
@@
Additive Sharing. To perform linear operations, additive sharing [38, 39, 40, 41] is a common MPC technique that splits private data into multiple secret shares, each distributed to a different party. In this approach, the data is secret-shared such that no single party can reconstruct it independently. If there are $j$ parties, a private value $X$ is represented as the sum of its shares:
$X=\sum[x]_{j}$
@@
    $[Z]_{j}=[X]_{j}+[Y]_{j},\quad Z=\sum\text{(locally computed)}\ [Z]_{j}$
  • Multiplication: To perform multiplication over additive shares, Beaver triples [42] are used. If $[Z]=[X]\cdot[Y]$, and the precomputed triple is $[c]=[a]\cdot[b]$ [19], each party $P_{j}$ does the following:
@@
      $Z_{j}=j\cdot E\cdot F+F\cdot a_{j}+E\cdot b_{j}+c_{j}$
Function Secret Sharing (FSS). Unlike additive sharing, Function Secret Sharing [43, 44, 45] splits the function itself across parties. That is, each party holds a share of the function rather than the data, and computation is done by applying these function shares to the input. If $f(X)$ is the function to evaluate and $X$ is the (possibly masked) input, then:
$f(X)=f_{1}(X)+f_{2}(X)$
Fig. 4(b) illustrates this: the private input $X$ is masked as $X_{c}=X+R$, and each party evaluates its function share on $X_{c}$. The final result is reconstructed by summing the outputs:
$f(X_{c})=f_{1}(X_{c})+f_{2}(X_{c})$
@@
4.4 AriaNN: Federated Learning Extension
Ryffel et al. [22] extended their secure training framework to support federated learning, enabling privacy-preserving model training across distributed clients. In this FL-based variant of AriaNN, the server is treated as an untrusted third party by both clients and model owners.
The training process begins by secret-sharing the model parameters between the server and clients, each holds one share of the model. As illustrated in Fig. 5, the protocol proceeds as follows: (1) Each client secret-shares a portion of its local data with the server while retaining the other share locally. (2) Clients and the server then collaboratively train their respective shares of the local model using MPC over a number of training iterations. (3) After local training, one designated client (e.g., $C_{A}$) acts as the aggregator, collecting secret-shared updates from other clients. (4) The aggregator performs secure weighted averaging of these updates, ensuring that individual data and model privacy are preserved throughout. The server, in parallel, aggregates its corresponding shares of the model updates. (5) The updated model share $[GM]_{1}$ is redistributed privately by the aggregator to all clients for the next round, while the server also updates its copy $[GM]_{0}$ accordingly.
@@ 4.4 AriaNN: Federated Learning Extension
5 Fully Secure Federated Learning (FuSeFL)
@@ 5.1 Overview of FuSeFL
    Model Distribution: The global model is secret-shared between two servers ($S_{1}$, $S_{2}$), who then transmit the shares to the two clients ($C_{k}$, $C_{l}$) in each group. Specifically, $S_{1}$ sends $[GM]_{1}$ to $C_{k}$, and $S_{2}$ sends $[GM]_{2}$ to $C_{l}$.
@@
    Secure Group Training: Clients $C_{k}$ and $C_{l}$ collaboratively train a local model over their secret-shared data ($D_{k}$ and $D_{L}$) using MPC. The model itself is maintained in a secret-shared form throughout training, ensuring confidentiality of both inputs and intermediate states.
@@
    Update Submission: After completing local training for several epochs, each client returns their share of the updated local model to the corresponding server—$C_{k}$ sends $[LM]_{1}$ to $S_{1}$ and $C_{l}$ sends $[LM]_{2}$ to $S_{2}$.
@@
5.2 Risk-Aware Dynamic Client Grouping
@@
s.t. each $\text{each }C_{i}\text{ appears exactly once.}$.
@@
Pipelined Execution: FuSeFL performs group assignment computation entirely in parallel with training. Because group matching relies solely on static metadata and does not interact with model state or local data, assignments for round $t{+}1$ are computed concurrently while training for round $t$ is ongoing. As a result, this procedure incurs no additional latency in the training workflow.
@@
5.3 Training Modes: Serial and Parallel
@@
In contrast, FuSeFL-Parallel (Fig. 7(b)) accelerates training by allowing both clients to train separate local models concurrently. The group maintains two model instances—$LM_{1}$ and $LM_{2}$—which are secret-shared among the clients. Each model is trained independently on $D_{1}$ and $D_{2}$ in parallel. After training, the clients locally aggregate the two updated models before securely transmitting the result to the servers. This mode reduces training latency and is better suited for time-sensitive deployments (evaluated in §7.2), at the cost of slightly higher memory usage and a modest trade-off in final model accuracy (evaluated in §7.1). Details are outlined in Algorithm 2).
@@
            Global model $[GM]_{j}$;
            Local models $[LM_{(i/2)}]_{j}$;
            Local datasets $[D_{i}]_{j}$
@@
                         Server s sends $[GM]_{j}$ to client c
@@
                  Client c $\leftarrow$ $[D_{c}]_{j}$
                   Client c $\leftarrow$ $[D_{(j-c)}]_{(j-c)}$
                   $[LM_{g}]_{j}$leftarrow $[GM]_{j}$
@@
                   clients train $[LM_{g}]$ using $[D_{0}]$
@@
                   clients train $[LM_{g}]$ using $[D_{1}]$
@@
                         client c sends $[LM_{g}]_{j}$ to server s
@@
                         Server s sends $[GM]_{j}$ to client c
@@
                  Client c $\leftarrow$ $[D_{c}]_{j}$
                   Client c $\leftarrow$ $[D_{(j-c)}]_{(j-c)}$
                   $[LM_{c}]_{j}$leftarrow $[GM]_{j}$
@@
                   clients train $[LM_{c}]$ using $[D_{0}]$
@@
                   clients train $[LM_{g}]$ using $[D_{1}]$

### v2

@@ Abstract
I Introduction
Federated Learning (FL) is a distributed machine learning paradigm that enables multiple clients to collaboratively train a global model without sharing their local data [1, 2, 3, 4, 5, 6, 7]. This paradigm is particularly suited for scenarios where legal, privacy, or infrastructural constraints prevent centralized data collection [8, 5]. In FL, a central server distributes a model to the clients, who perform local training and return updates, such as gradients or model weights. These updates are aggregated to refine the global model, all while the raw data remains on client devices, preserving confidentiality and reducing transmission costs. While FL offers clear advantages in preserving data locality, it remains vulnerable to indirect privacy breaches. In particular, submitted gradients or model updates can unintentionally encode patterns from the underlying data, enabling adversaries to extract sensitive information from these values [9, 4, 10, 11].
To mitigate such risks, a wide range of techniques have been proposed to protect the privacy of clients’ data during training. Homomorphic encryption (HE) enables secure aggregation over encrypted updates without decryption [3, 4], but incurs high computational and memory costs, often slowing down training by several orders of magnitude [12]. Differential privacy (DP) defends against inference attacks by adding noise to updates [5, 6], though stronger privacy budgets degrade model accuracy [12]. Secure multi-party computation (MPC) allows collaborative aggregation without exposing individual inputs [7, 13, 11, 14, 4, 15], but suffers from growing communication overhead as the number of clients increases [12].
Beyond data privacy, a critical yet often underappreciated concern in federated learning is the privacy of the model itself, particularly from the perspective of the model owner. In typical FL deployments, the global model, often developed through significant intellectual and financial investment, is shared with all participating clients, some of whom may be semi-honest or untrusted [16]. This exposes the model to risks such as extraction, reverse engineering, or unauthorized use. The concern becomes even more pressing in cross-silo FL, where clients are typically large institutions, such as hospitals, banks, or research labs, that operate under strict privacy regulations and contractual obligations. In such settings, the model is not only a valuable proprietary asset but also a potential vector for secondary leakage of sensitive information. For example, attackers may perform model inversion to reconstruct private inputs or launch membership inference attacks to determine whether a particular data point was part of the training set [17]. Ensuring model confidentiality is therefore essential, not only to protect intellectual property but also to comply with legal, regulatory, and contractual boundaries.
@@ I Introduction
WW-FL [20] adopts a fully offloaded training architecture in which secure multi-party computation (MPC) is performed across distributed clusters of servers. While this design offers strong security guarantees, it requires clients to upload their entire datasets to third-party servers, contradicting the foundational principle of cross-silo FL, where data should remain local to each silo. In addition, this architecture introduces a trade-off between communication overhead, memory usage, and the number of server clusters. When dealing with thousands of clients, increasing the number of servers raises the system’s cost while reducing communication time and memory usage. However, reducing the number of server clusters leads to significant communication overhead, as large volumes of sensitive data must be transmitted over the network, increasing both latency and memory consumption. Moreover, WW-FL depends on geographically distributed compute infrastructure, which elevates deployment complexity and operational costs. In cross-silo environments, where clients such as hospitals or financial institutions already possess rich compute and storage capabilities, outsourcing training to external servers is not only inefficient but also unnecessary.
@@
AriaNN [25] introduces a hybrid federated scheme that jointly protects model and data privacy by secret-sharing the model between the clients and server and distributing the computation using MPC. This design prevents either party from observing the full model or training data in plaintext. While more aligned with FL principles, AriaNN suffers from serious scalability limitations. Each client must train its share of the model in coordination with the server, typically in a sequential fashion due to server-side compute constraints. Communication overhead grows linearly with client count; even modest datasets like MNIST can require tens of megabytes per client, and this cost becomes prohibitive for large datasets common in siloed environments. Furthermore, the server must maintain a dedicated model instance per client in addition to the global model, leading to unsustainable memory overhead for large-scale models such as VGG-16.
@@
In FuSeFL, clients are dynamically paired to form collaborative groups. Within each group, the global model is secret-shared between the two clients, and training is jointly executed using secure multi-party computation (MPC) over their secret-shared local data. This structure prevents any party from accessing the model or data in plaintext. After training, each group submits secret-shared updates to two non-colluding aggregation servers, which perform secure aggregation without learning individual contributions.
@@
  • A fully secure FL scheme: We develop FuSeFL, a federated learning scheme that holistically secures model parameters, client data, and intermediate updates using lightweight MPC across client groups, addressing both model and data privacy within a unified design (§V-A).
  • Decentralized and scalable training: FuSeFL eliminates server-side training bottlenecks by distributing computation among client pairs. This design supports parallelism, avoids reliance on third-party compute infrastructure, and is well-suited for cross-silo FL environments with powerful clients (§VII-B).
  • Practical efficiency with robust privacy guarantees: Our approach significantly reduces memory and communication overheads while mitigating key inference threats such as model inversion, gradient leakage, and membership inference, making FuSeFL scalable to real-world deployments involving many silos (§VII-D, §VII-C).
II Motivation and Goal
II-A Limitations of Existing Work
@@
    Model Exposure and Trust Assumptions: Many existing FL schemes expose the full global model to clients during training or inference, risking model theft or misuse, particularly in regulated domains where model intellectual property is sensitive [28, 18, 19]. Moreover, several privacy-preserving methods, such as WW-FL[20], offload client data or models to trusted third-party servers using MPC or HE protocols [22, 23, 24, 25, 29, 30]. These trust assumptions violate the core principle of cross-silo FL, which requires data to remain within institutional boundaries [8, 31].
@@
    Communication and Storage Overheads: Secure training introduces high communication and memory demands. For example, training Network1 over MNIST with AriaNN-FL requires more than 10 MB of data exchange per client per epoch [25]. With hundreds of clients, this quickly overwhelms bandwidth limits. Additionally, the server must retain a copy of each client’s model update until aggregation is complete. For large models like VGG-16, this results in excessive memory usage: as illustrated in Fig. 1 (right axis), AriaNN-FL stores 128 full model instances when training with 128 clients. In schemes such as WW-FL, where the model and user data are offloaded to third-party servers, communication latency becomes significant when submitting user data to the training server cluster, as the number of server clusters is limited due to their associated cost.
@@ II-B Our Goals
III Threat Model and Security Considerations
@@
These guarantees mitigate risks such as unauthorized observation, gradient leakage, and reconstruction attacks. Since the model is never exposed in plaintext outside the trusted boundary, no untrusted party can access it to carry out membership inference or model inversion attacks aimed at extracting information about the training user data. FuSeFL enforces these protections using efficient secure MPC protocols.
@@
IV Background
IV-A Notations and Preliminaries
Our scheme leverages MPC over the ring $\mathbb{Z}_{2^{n}}$ to ensure the confidentiality of training data and model throughout the computation. Table I provides a summary of the key notations used in the remainder of the paper.
@@ IV-A Notations and Preliminaries
GMGM Global model in plaintext
$GM]_{j}$ $j^{th}$ share of the secret shared global model
@@
$LM_{i}]_{j}$ $j^{th}$ share of the secret shared $i^{th}$ local model
@@
$D_{i}]_{j}$ $j^{th}$ share of the secret shared data of $i^{th}$ client
IV-B Federated Learning (FL)
@@
IV-C Secure Multi-Party Computation (MPC)
@@
Additive Sharing. To perform linear operations, additive sharing [38, 39, 40, 41] is a common MPC technique that splits private data into multiple secret shares, each distributed to a different party. In this approach, the data is secret-shared such that no single party can reconstruct it independently. We define additive secret sharing over a finite ring $\mathbb{Z}_{2^{n}}$, where $n$ is the bit-length of encoded values and arithmetic is performed modulo $2^{n}$.
Let $X\in\mathbb{Z}_{2^{n}}$ be a private value. In a setting with j parties, $X$ is split into j additive shares ($x]_{j}$) and is represented as the sum of its shares:
$X=\sum[x]_{j}\mod{2^{n}}$
@@
    $Z]_{j}=[X]_{j}+[Y]_{j}\mod{2^{n}}$
    $\quad Z=\sum\text{(locally computed)}\ [Z]_{j}\mod{2^{n}}$
  • Multiplication: To perform multiplication over additive shares, Beaver triples [42] are used. If $Z]=[X]\cdot[Y]$, and the precomputed triple is $c]=[a]\cdot[b] [$22], each party $P_{j}$ does the following:
@@
      $Z_{j}=j\cdot E\cdot F+F\cdot a_{j}+E\cdot b_{j}+c_{j}\mod{2^{n}}$
Function Secret Sharing (FSS). Unlike additive sharing, Function Secret Sharing [43, 44, 45] splits the function $f:\mathbb{Z}_{2^{n}}\rightarrow\mathbb{Z}_{2^{n}}$ across parties. Each party holds a share of the function rather than the data, and computation is done by applying these function shares to the input.
If $f(x)$ is the function to evaluate and $x$ is the (possibly masked) input, then:
$f(x)=f_{1}(x)+f_{2}(x)\mod{2^{n}}$
Fig. 4(b) illustrates that the private input $X$ is masked as $X_{c}=X+R$, and each party evaluates its function share on $X_{c}$. The final result is reconstructed by summing the outputs.
@@
IV-D AriaNN: Federated Learning Extension
Ryffel et al. [25] proposed AriaNN, which uses additive sharing and function secret sharing (FSS) [43, 45] to privately perform deep learning tasks (both training and inference) with minimal communication overhead. In their scheme, both the data and the model are secret-shared between two parties to ensure that no single party can reconstruct the sensitive information. AriaNN supports several operations required for neural network training and inference.
AriaNN uses additive secret sharing to securely perform linear computations (explained in §IV-C). For example, to support secure multiplication and convolution, AriaNN employs Beaver triples [42], specifically matrix-based Beaver triples [22], for efficient batch operations.
To perform comparison-based computations like ReLU, AriaNN uses FSS [43, 45] (explained in §IV-C). By using FSS, AriaNN achieves a significant reduction in the number of communication rounds compared to previous works [40, 23, 24]. To implement comparison $(a\leq\alpha)$, AriaNN uses private equality testing proposed by [45]. The evaluators evaluate all possible paths before the special path (where the public value $a$ is equal to the private value $\alpha$). Then, they sum them up to obtain the final output of the comparison [25].
AriaNN, then, extended their secure training framework to support federated learning, enabling privacy-preserving model training across several distributed clients.
The training process begins by secret-sharing the model parameters between the server and clients, each holds one share of the model. As illustrated in Fig. 5, the protocol proceeds as follows: (1) Each client secret-shares a portion of its local data with the server while retaining the other share locally. (2) Clients and the server then collaboratively train their respective shares of the local model using MPC over a number of training iterations. (3) After local training, one designated client (e.g., $C_{A}$) acts as the aggregator, collecting secret-shared updates from other clients. (4) The aggregator performs secure weighted averaging of these updates, ensuring that individual data and model privacy are preserved throughout. The server, in parallel, aggregates its corresponding shares of the model updates. (5) The updated model share $GM]_{1}$ is redistributed privately by the aggregator to all clients for the next round, while the server also updates its copy $GM]_{0}$ accordingly.
@@ IV-D AriaNN: Federated Learning Extension
V Fully Secure Federated Learning (FuSeFL)
@@ V-A Overview of FuSeFL
    Model Distribution: The global model is secret-shared between two servers ($S_{1}$, $S_{2}$), who then transmit the shares to the two clients ($C_{k}$, $C_{l}$) in each group. Specifically, $S_{1}$ sends $GM]_{1}$ to $C_{k}$, and $S_{2}$ sends $GM]_{2}$ to $C_{l}$.
@@
    Secure Group Training: Clients $C_{k}$ and $C_{l}$ collaboratively train a local model over their secret-shared data ($D_{k}$ and $D_{L}$) using MPC. In our case, we use AriaNN to perform secure training within a group (§IV-D). The model itself and the data are maintained in a secret-shared form throughout training, ensuring confidentiality of both inputs and intermediate states. One share of the data remains local on each client to ensure that the data remains confidential and never leaves the client in its entirety.
@@
[…]
@@
V-B Risk-Aware Dynamic Client Grouping
@@
s.t. $\text{each }C_{i}\text{ appears exactly once.}$
@@
Pipelined Execution: FuSeFL performs group assignment computation entirely in parallel with training. Because group matching relies solely on static metadata and does not interact with model state or local data, assignments for round t+1t{+}1 are computed concurrently while training for round $t$ is ongoing. As a result, this procedure incurs no additional latency in the training workflow.
@@
V-C Training Modes: Serial and Parallel
@@
In contrast, FuSeFL-Parallel (Fig. 7(b)) accelerates training by allowing both clients to train separate local models concurrently. The group maintains two model instances ($LM_{1}$ and $LM_{2}$), which are secret-shared among the clients. Each model is trained independently on $D_{1}$ and $D_{2}$ in parallel. After training, the clients locally aggregate the two updated models before securely transmitting the result to the servers. This mode reduces training latency and is better suited for time-sensitive deployments (evaluated in §VII-B), at the cost of slightly higher memory usage and a modest trade-off in final model accuracy (evaluated in §VII-A). Details are outlined in Algorithm 2).
@@
     Global model $GM]_{j}$;
     Local models $LM_{(i/2)}]_{j}$;
     Local datasets $D_{i}]_{j}$
@@
           Server s sends $GM]_{j}$ to client c
@@
       Client c $\leftarrow$ $D_{c}]_{j}$
        Client c $\leftarrow$ $D_{(j-c)}]_{(j-c)}$
        $LM_{g}]_{j} \leftarrow [$$GM]_{j}$
@@
        clients train $LM_{g}]$ using $D_{0}]$
@@
        clients train $LM_{g}]$ using $D_{1}]$
@@
           client c sends $LM_{g}]_{j}$ to server s
@@
           Server s sends $GM]_{j}$ to client c
@@
       Client c $\leftarrow$ $D_{c}]_{j}$
        Client c $\leftarrow$ $D_{(j-c)}]_{(j-c)}$
        $LM_{c}]_{j} \leftarrow [$$GM]_{j}$
@@
        clients train $LM_{c}]$ using $D_{0}]$
@@
        clients train $LM_{g}]$ using $D_{1}]$